Paul is 15 years younger then Ken. In 5 years Ken will be twice as old as Paul. How old will Paul be in four years?
Answer by ewatrrr(24785)


Hi
Paul is 15 years younger then Ken
Let x and (x-15) represent Ken and Paul's ages NOW respectively
Write as we Read*** in five years...
[x+5] = 2[(x-15) + 5]
solving for x
x + 5 = 2x - 20
25 = x, Ken's age now, Paul is now 10. Paul be in four years will be 14.
CHECKING our Answer***in five years...
30 = 2*15